# Quantitative entropy estimates for 2D stochastic vortex model on the whole space under moderate interactions

### 05 Mar 2026 v3

@@ 2.1 Time evolution of the relative entropy
$II_{t}+IV_{t}$ $=\frac{1}{2N^{2}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{1}{\rho_{s}^{N}}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,ds\,dx,$ (31)
$III_{t}$ $=-\frac{1}{2}\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds,$ (32)
@@
$M_{t}^{N}$ $=\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dB_{s}\,dx$
$+\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
$-\frac{1}{N}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\int_{0}^{t}\nabla V^{N}(x-X_{s}^{i})\,dW^{i}_{s}\,dx$
$+\int_{\mathbb{R}^{d}}\int_{0}^{t}\frac{\rho_{s}^{N}}{\rho_{s}}\sigma^{\top}_{s}\nabla\rho_{s}\,dB_{s}\,dx.$ (33)
2.2 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
@@
By decay of $\rho_{0}$ in (12), along with Borel-Cantelli Lemma, almost surely, there exists $N_{0}=N_{0}(\omega)$ such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
@@
With this in view for all $\omega\in\Lambda_{0}$, with
$\Lambda_{0}\doteq\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}$ (40)
there exists $N_{0}=N_{0}(\omega)$, such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$, since $t\mapsto X_{t}^{i,N}(\omega)$ is continuous, almost surely.
@@
Now, if $|x|\leq 2|X_{s}^{i,N}|$, then $|x|\leq 2N^{\beta}$ for $s\in(0,\tau^{N})$, and since $V^{N}\leq N^{\beta}$, we derive for $\alpha>d/2$
@@
$\phi(x)\doteq\left(\left(1+|x|^{2}\right)^{-\alpha}+\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right)$ (45)
@@
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
$\overset{\alpha>d/2}{\leq}C_{\alpha,T}N^{-\theta_{1}},$ (46)
@@
2.3 Estimates for $I_{t}$: the nonlinear terms
@@
$\doteq I^{1}_{t}-I^{2}_{t},$ (47)
where
$\doteq-I^{1,1}_{t}-I^{1,2}_{t}.$ (48)
$\leq\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{2}}{\rho_{s}^{N}}\left|\nabla\left(\frac{\rho_{s}^{N}}{\rho_{s}}\right)\right|^{2}\,dx\,ds+C_{\epsilon}\int_{0}^{t}\|K_{0}\|_{\infty}^{2}\|\rho_{s}-\rho_{s}^{N}\|_{1}^{2}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{s}|^{2}}{\rho_{s}^{2}}\rho_{s}^{N}\,dx\,ds.$ (49)
We observe that, for $\eta>2C_{1}(8T+C_{3})+1$
which implies by Donsker-Varadhan inequality (108),
@@
Hence by Csiszár-Kullback-Pinsker inequality (107) and (49), we derive
Regarding term $I_{t}^{1,1}$ in (48), by convolution inequality, we find
$I_{t}^{1,1}$ $=\int_{0}^{t}\|K_{0}\|_{\infty}\|\rho^{N}_{s}-\rho_{s}\|_{1}\int_{\mathbb{R}^{d}}\left|\frac{\nabla^{2}\rho_{s}}{\rho_{s}}\right|\rho^{N}_{s}\,dx\,ds.$ (53)
In addition, if $C_{1,2}\doteq\max{(C_{1},C_{2})}$ since
@@
for $\eta>2C_{1,2}(8T+C_{3})+1$, we obtain
It follows by Donsker-Varadhan inequality (108),
@@
Therefore, by (56) in (53), we get
@@
$\leq\int_{\mathbb{R}^{d}}2\|K_{0}\|_{\infty}|\nabla\rho_{s}^{N}(y)|\,dy$
$=2\|K_{0}\|_{\infty}\int_{\mathbb{R}^{d}}|\nabla\rho_{s}^{N}(y)|\,dy.$ (59)
Now from Leibniz’s rule we deduce,
@@
which implies, by $\epsilon$-Young and Holder inequalities,
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(1/2)(\rho^{N}_{s})^{1/2}(\rho^{N}_{s})^{1/2}\left(C_{\eta}\right)^{\frac{1}{2}}\,dx\,ds$
$\overset{\text{Young}}{\leq}\epsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{N}_{s}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|^{2}\,dx\,ds+C_{\epsilon,\eta}\int_{0}^{t}\ \mathcal{H}(\rho^{N}_{s}|\rho_{s})\,ds$
[…]
2.4 Application of Gronwall’s Lemma
In this subsection, we conclude our localizated estimates, by applying Gronwall’s inequality.
Indeed, with the same computations as in [40], for all $\delta\in(0,\theta_{2})$ there exists a random variable $A_{0}$, finite almost surely, such that
$\sup_{t\in[0,T]}|M_{t}^{N}|\leq A_{0}N^{-\theta_{2}+\delta},$ (64)
with $\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)$.
[…]
$-\int_{0}^{t\wedge\tau^{N}}\mathcal{I}(\rho_{s}^{N}|\rho_{s})\,ds$
$+C_{\alpha,T}N^{-\theta_{1}}+A_{0}N^{-\theta_{2}+\delta}+TC_{\eta}$ (65)
and then for $\eta>>1$ and $\epsilon<<1$, we deduce
Finally, from Gronwall’s inequality (109) in (66), recalling (36), (39) and (40), for all $\omega\in\Lambda_{0}$, there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, $\tau^{N}(\omega)>0$ and
@@ 2.4 Application of Gronwall’s Lemma
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t\wedge\tau^{N}}^{N}|\rho_{t\wedge\tau^{N}})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right)$ (67)
with
2.5 End of the proof: Removing the stopping time
$=\left\{\exists i;\sup_{t\in[0,T]}|X_{t}^{i,N}|\geq N^{\beta}\right\}$
$\subset L_{1}^{N}\cup L_{2}^{N}\cup L_{3}^{N},$ (69)
with
and
We note that, with the same computations like those led (37), we deduce
We now take care about the $L_{2}^{N}$. Indeed, we note that $W_{t}^{i,N}=\int_{0}^{t}dW_{s}^{i,N}$, and since
$|W_{t}^{i,N}|^{2}=\sum_{j=1}^{d}(W_{t}^{i,N,j})^{2},$
if $|W_{t}^{i,N}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}.$ Therefore,
$\left\{\sup_{t\leq T}|W_{t}^{i,N}|\geq a\right\}\subset\bigcup_{j=1}^{d}\left\{\sup_{t\leq T}|W_{t}^{i,N,j}|\geq\frac{a}{\sqrt{d}}\right\},$ (74)
we obtain
[…]
$\leq\sqrt{\frac{T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2T)\right)},$ (75)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [26].
Now we will derive an analogous estimate for $L_{3}^{N}$. First, we observe that, if $\sigma=0$ we are done, thus suppose $\sigma\neq 0$. Since $\sigma$ is bounded, for $j\in\{1,...,d\}$, the function defined by $t\mapsto\langle X^{j}\rangle_{t}$ is continuous and increasing. Hence, recalling (19), by Theorem 4.6, in Chapter 3 of [26], for all $j\in\{1,...,d\}$, we get
$\sup_{t\leq T}|X_{t}^{j}|=\sup_{t\leq T}\left|\widetilde{B}^{j}_{\langle X^{j}\rangle_{t}}\right|\leq\sup_{t\leq\|\sigma\|_{\infty}^{2}T}|\widetilde{B}^{j}_{t}|,$ (76)
where $(\widetilde{B}^{j}_{t})_{t\geq 0}$ is the time-change process, which in particular, is a one dimensional standard Brownian motion. In addition, since
$|X_{t}|^{2}=\sum_{j=1}^{d}|X_{t}^{j}|^{2}$
if $|X_{t}|\geq a$, then there exists j=1,,dj=1,...,d, such that $|X_{t}^{j}|\geq\frac{a}{\sqrt{d}}.$ Therefore, for all $a>0$, we find
[…]
It follows that, by (76) and (77)
[…]
$\leq\sqrt{\frac{\|\sigma\|_{\infty}^{2}T}{2\pi}}\frac{4dN}{\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}}$
$\times\exp{\left(-\left(\frac{C_{K_{0},V}N^{\beta}}{\sqrt{d}}\right)^{2}/(2\|\sigma\|_{\infty}^{2}T)\right)},$ (78)
where in the second inequality, we are using (8.3)’ in Chapter 2 of [26].
Hence, by (73), (75) and (78) in (69), we find
$\sum_{N\in\mathbb{N}}\mathbb{P}\left(\tau^{N}<T\right)<\infty$ (79)
and then by Borel-Cantelli Lemma, we arrive at
Finally, by (40) and (80), for $T<\min{\left(T_{1};(1/\|\nabla V\|_{1}\|K_{0}\|_{\infty})\right)}$ for all $\omega\in\Lambda$, with
$\Lambda\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}\{\tau^{N}=T\}\right)\cap\left\{A_{0}<\infty\right\},$ (81)
there exists $N_{0}(\omega)$ such that for $N\geq N_{0}(\omega)$, we have $\tau^{N}(\omega)=T$ and
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (82)
In particular, by (20) and (64), we conclude with
$\limsup_{N\to\infty}\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim T,$ (83)
almost surely.
3 Proof of Theorem 2
In this section we present the proofs of Theorem 2. In Subsection 3.1, by applying the Itô’s formula, such as in [27], we derive an evolution equation for the energy of the difference between regularized empirical measure and the solution of (1). In Subsection 3.2, we address the nonlinear terms, combining bounds on the Fisher information of the particle system with the Ladyzhenskaya inequality.
Finally, in Subsections 3.3, combining the previous estimates with a nonlinear Gronwall’s lemma, we complete the proof.

### 13 Apr 2026 v4

@@ 2.1 Time evolution of the relative entropy
$II_{t}+IV_{t}$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{1}{\rho_{s}^{N}}|\nabla V^{N}(x-X_{s}^{i,N})|^{2}\,dx\,ds,$ (28)
$III_{t}$ $=-\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds,$ (29)
@@
$M_{t}^{N}$ $=\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dx\,dB_{s}$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\sigma^{\top}_{s}\nabla\rho_{s}^{N}\,dx\,dB_{s}$
$+\frac{1}{N}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}\left[\ln(\rho_{s})-\ln(\rho_{s}^{N})\right]\nabla V^{N}(x-X_{s}^{i})\,dx\,dW^{i,N}_{s}$
$-\frac{1}{N}\sum_{i=1}^{N}\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla V^{N}(x-X_{s}^{i})\,dx\,dW^{i,N}_{s}$
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\frac{\rho_{s}^{N}}{\rho_{s}}\sigma^{\top}_{s}\nabla\rho_{s}\,dx\,dB_{s}.$ (30)
2.1.1 Estimates for $II_{t}+IV_{t}$: quadratic variation terms
@@
By decay of $\rho_{0}$ in (11), along with Borel-Cantelli Lemma, there exists $\Lambda_{0}\in\mathcal{F}$, $\mathbb{P}(\Lambda_{0})=1$, with the following property: for all $\omega\in\Lambda_{0}$, there exists $N_{0}=N_{0}(\omega)$ such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
@@
With this in view, since $t\mapsto X_{t}^{i,N}$ is continuous, for all $i\in\{1,...,N\}$, almost surely, by setting
$\Lambda_{X_{0}}\doteq\bigcap_{N\in\mathbb{N}}\bigcap_{i=1}^{N}\left\{t\mapsto X_{t}^{i,N}\,\,\,\text{is continuous}\right\}$
and
$\Lambda_{0}\doteq\left(\bigcup_{M=1}^{\infty}\bigcap_{N=M}^{\infty}(\Lambda^{N})^{c}\right)\cap\Lambda_{X_{0}}$ (37)
we have that $\mathbb{P}(\Lambda_{0})=1$, and for all $\omega\in\Lambda_{0}$, there exists $N_{0}=N_{0}(\omega)$, such that $\tau^{N}(\omega)>0$, for all $N\geq N_{0}$.
@@
Now, if $|x|\leq 2|X_{s}^{i,N}|$, then $|x|<2N^{\beta}$ for $s\in(0,\tau^{N})$, and since $V^{N}\leq N^{\beta}$, we derive for $\alpha>d/2$
@@
$\phi(x)\doteq\left(\left(1+|x|^{2}\right)^{-\alpha}+\left(1+\frac{|x|^{2}}{4}\right)^{-\alpha}\right).$ (42)
@@
$\overset{}{\leq}C_{\alpha,d}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N^{2}}\int_{0}^{t\wedge\tau^{N}}\sum_{i=1}^{N}\int_{\mathbb{R}^{d}}\frac{|V^{N}(x-X_{s}^{i,N})|}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha,d}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\frac{\frac{1}{N}\sum_{i=1}^{N}V^{N}(x-X_{s}^{i,N})}{\rho^{N}_{s}(x)}\phi(x)\,dx\,ds$
$\overset{}{\leq}C_{\alpha,d}\frac{N^{\beta(1+\frac{2}{d}+2\alpha)}}{N}\int_{0}^{t\wedge\tau^{N}}\int_{\mathbb{R}^{d}}\phi(x)\,dx\,ds$
$\overset{\alpha>d/2}{\leq}C_{\alpha,d,T}N^{-\theta_{1}},$ (43)
@@
2.1.2 Estimates for $I_{t}$: the nonlinearity
We now address the nonlinear terms in (26). The main novelty is to use the stopping time obtained in previous Subsection, along with the dissipation that comes from of Fisher information and Ladyzenkaya inequality, to tackle with the singularity of kernel $K$.
@@ 2.1.2 Estimates for $I_{t}$: the nonlinearity
$-\int_{0}^{t}\int_{\mathbb{R}^{d}}\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)]\right\rangle\,dx\,ds$
$\doteq I_{t}^{1}+I_{t}^{2}$ (44)
Thus, by Young inequality together with (7), we note that
$I_{t}^{1}$ $\leq\epsilon\int_{0}^{t}\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)\,ds+C_{\epsilon}\int_{0}^{t}\left(\|\rho_{s}^{N}-\rho_{s}\|_{1}^{2}+\|\rho_{s}^{N}-\rho_{s}\|_{4}^{2}\right)\,ds$
$\leq\epsilon\int_{0}^{t}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+\|\nabla(\rho^{N}_{s}-\rho_{s})\|_{2}^{2}\right)\,ds$
$\hskip 80.29999pt+C_{\epsilon,d}\int_{0}^{t}\left(\|\rho_{s}^{N}-\rho_{s}\|_{1}^{2}+\|\rho_{s}^{N}-\rho_{s}\|_{2}^{2}\right)\,ds,$ (45)
where, in the second step, we use the Ladyzhenskaya and Young inequalities.
Now, we will focus on the term $I_{t}^{2}$ in (44). First, we observe that by subtracting and adding $K\ast\rho_{s}(\cdot)$ and $K\ast\rho_{s}(x)$ along with (7)
$\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|$ $\lesssim\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}+\|\rho^{N}_{s}-\rho_{s}\|_{4}\right)$
$\hskip 80.29999pt+\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|.$ (46)
Additionally, by (7) and (17), we obtain
$\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\right\rangle$
@@
We define
$R\doteq\left\{(x,\cdot);|x-\cdot|<N^{-\beta\gamma/d}\right\}$ (48)
It follows by (47) and (48) that
$\int_{\mathbb{R}^{d}}\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\mathds{1}_{R}(x,\cdot)\right\rangle\lesssim N^{-\beta\gamma^{2}/d}.$ (49)
On the other hand, by decay of $V$, given by (6), we have
$V^{N}(x-\cdot)|x-\cdot|^{\gamma}$ $=N^{\beta}V(N^{\beta/d}(x-\cdot))|x-\cdot|^{\gamma}$
@@
$\lesssim N^{-3+\beta}\left|(x-\cdot)\right|^{-3d/\beta}.$ (50)
Hence, we derive
$\int_{\mathbb{R}^{d}}\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\,dx$
@@
$\hskip-170.6375pt+N^{-3+\beta}\int_{N^{\beta}<|x|\leq 2N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx$
$\hskip-170.6375pt+N^{-3+\beta}\int_{|x|>2N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx$
$\hskip-90.3375pt\doteq I^{1}+I^{2}+I^{3}.$ (51)
Now, for $\gamma\in(0,1/2)$, we get
$I^{1}$ $=N^{-3+\beta}\int_{|x|\leq N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx\leq N^{-3+\beta}N^{d\beta}N^{3\gamma}$
$\leq N^{-3+(1+d)\beta+3\gamma}.$ (52)
Also, for the same reason we arrive at,
$I^{2}$ $=N^{-3+\beta}\int_{N^{\beta}<|x|\leq 2N^{\beta}}\left|\left\langle S_{s}^{N},\left|(x-\cdot)\right|^{-3d/\beta}\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\right|\,dx\leq C_{d}N^{-3+\beta}N^{d\beta}N^{3\gamma}$
$\leq C_{d}N^{-3+(1+d)\beta+3\gamma}.$ (53)
Moreover, if $|x|>2N^{\beta}$, since that for $s\in(0,\tau^{N})$, $|X_{s}^{i,N}|<N^{\beta}$ (see (33)), for all $i\in\{1,...,N\}$, we have, $|x-\cdot|\geq 1/2|x|$, which implies
$\int_{|x|>2N^{\beta}}|x-\cdot|^{-3d/\beta}\,dx\leq C_{d}\int_{|x|>2N^{\beta}}|x|^{-3d/\beta}\,dx=C_{d}N^{-3d+d\beta}$
and then,
$I^{3}$ $=N^{-3+\beta}\int_{\mathbb{R}^{d}}\int_{|x|>2N^{\beta}}|x-y|^{-3d/\beta}\mathds{1}_{R^{c}}(x,y)\,dx\,S_{s}^{N}(dy)$
$\leq N^{-3+\beta}C_{d}(N^{\beta})^{-3d/\beta+d}\leq C_{d}N^{-3+(1+d)\beta+3\gamma}.$ (54)
It follows by (51)-(54) that,
$\int_{\mathbb{R}^{d}}\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)\right|\mathds{1}_{R^{c}}(x,\cdot)\right\rangle\lesssim N^{-3+(1+d)\beta+3\gamma}.$ (55)
Therefore, by (49) and (55), $\theta_{2}\doteq\beta\gamma^{2}/d$ with $\gamma\in(0,1/2)$, and $s\in(0,\tau^{N})$, we deduce,
$\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|\left\langle S_{s}^{N},V^{N}(x-\cdot)\left|K\ast\rho_{s}^{N}(\cdot)-K\ast\rho_{s}^{N}(x)\right|\right\rangle\,dx$
@@
$+\int_{\mathbb{R}^{d}}\left|\nabla\ln\left(\frac{\rho_{s}}{\rho_{s}^{N}}\right)\right|(\rho_{s}^{N})^{1/2}\left|\left\langle S_{s}^{N},V^{N}(x-\cdot)[K\ast\rho_{s}(\cdot)-K\ast\rho_{s}(x)]\right\rangle\right|^{1/2}\,dx$
@@
$\hskip 120.45pt+C_{\epsilon,\gamma,\beta,d}\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}^{2}+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}+N^{-\theta_{2}}\right),$ (56)
where the last step follows from the Ladyzhenskaya and Young inequalities.
Finally, by (56), we get
$I_{t\wedge\tau^{N}}^{2}$ $\leq\int_{0}^{t\wedge\tau^{N}}\epsilon\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+\|\nabla(\rho_{s}^{N}-\rho_{s})\|_{2}^{2}\right)\,ds$
@@
$\hskip 80.29999pt+C_{\epsilon,\gamma,\beta,d}\int_{0}^{t\wedge\tau^{N}}\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}^{2}+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}+N^{-\theta_{2}}\right)\,ds,$ (57)
and thus, by (44), (45) and (57), we end up with
$I_{t\wedge\tau^{N}}$ $\leq 2\epsilon\int_{0}^{t\wedge\tau^{N}}\left(\mathcal{I}\left(\rho_{s}^{N}|\rho_{s}\right)+\|\nabla(\rho_{s}^{N}-\rho_{s})\|_{2}^{2}\right)\,ds$
$\hskip 80.29999pt+C_{\epsilon,\gamma,\beta,d}\int_{0}^{t\wedge\tau^{N}}\left(\|\rho^{N}_{s}-\rho_{s}\|_{1}^{2}+\|\rho^{N}_{s}-\rho_{s}\|_{2}^{2}+N^{-\theta_{2}}\right)\,ds,$ (58)
where $\theta_{2}\doteq\beta\gamma^{2}/d$ and $\gamma\in(0,1/2)$.